# From many-body ab initio to effective excitonic models: a versatile mapping approach including environmental embedding effects

### v1

@@ 1 Introduction
The accurate first-principles description of elementary excitations (excitons) in organic materials represents an invaluable investigation tool for understanding their photophysical properties as well as phenomena underlying applications in optoelectronics, such as energy transfer, exciton splitting or charge separation in organic solar cells.1, 2, 3, 4
@@
Time-dependent density-functional theory (TD-DFT), in its linear-response formulation within the adiabatic approximation,5 has been the most popular investigation method for the characterization of excitons in molecular systems. TD-DFT arguably represents one of the most reliable methods among those being computationally affordable for large and complex real-life materials. However, TD-DFT is known to present some limitations, such as the inaccurate and strongly functional-dependent description of CT excitations. This results in a possibly incorrect state ordering with FEs,6, 7 or in missing the characteristic Coulomb-like dependence of CT states energy with inter-molecular distance when using local, semilocal, or global hybrid functionals.8 These shortcomings can be significantly attenuated with the use of range-separated functionals,9, 10 leading to results that however depend on the the specific parameterization for the range separation of the Coulomb repulsion term.
@@
Many-body perturbation methods based on Green’s function theory, such as the 11, 12, 13, 14, 15, 16 and the Bethe-Salpeter equation (BSE) formalisms,17, 18, 19, 20 stand as robust and affordable options for computing the excited states in molecular systems.21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31 /BSE features the same computational scaling with respect to the system size as TD-DFT and provides an accurate description for excitations of different nature thanks to a proper description of electronic correlations. Benchmark studies against reference methods and experimental data demonstrated a typical accuracy of 0.1-0.2 eV for charged and optical excitations.32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43 Specifically, the explicit accounting for the non-local excitonic electron-hole interaction permits a faithful description of excitations regardless of their nature (FE vs. CT), of the type of system under scrutiny (e.g. organic homo- and hetero-molecular, inorganic, hybrid) and, importantly, of the functional used as starting point for the many-body perturbation theory.31, 36, 41
@@
The possibility of describing molecular excitations in condensed phases with Green’s function many-body method has been enabled by the recent development of a multiscale embedding method of different levels of detail, namely polarizable continuum model (PCM),44, 45, 46, 47 classical atomistic (polarizable QM/MM)48, 49, 30 and full-quantum (QM/QM’)50, 51. Embedded and BSE formalisms have been successfully applied in the contexts of molecular doping52, 53, 54, 55 and organic photovoltaics.56 Many-body perturbation theory methods have been historically developed for extended (i.e. periodic) systems14 and applications to molecular solids have been reported in the literature.57, 58, 59, 60 These calculations are extremely expensive and hardly affordable for realistic molecular solids encompassing many atoms in the unit cell or featuring structural disorder.
@@
Effective model Hamiltonians for excited states represent a cheap and insightful alternative to a full ab initio treatment for the description of the low-energy physics of multichromophoric systems.61, 3, 62, 63 These approaches rely on the definition of a reduced set of diabatic states characterized by wave functions localized on single molecular units, usually including pure intra-molecular FE and pure inter-molecular CT excitations with electron and hole localized on different fragments (see Figure 1). For this reason, these approaches are sometimes referred to as “low-energy” or “few-state” models. A typical model Hamiltonian for a molecular solid can be expressed as the sum of three terms,
@@
describing FEs (standard Frenkel exciton model), inter-molecular CT states, and the interaction between the two types of exciton (see SI for explicit expressions). Black arrows in Fig. 1 show the elementary processes corresponding to the quantum coupling between states, namely FE transfer(XT), excited-state hole/electron transfer (XHT/XET), and hole/electron transfer (HT/ET). The energy of the diabatic basis states and the coupling between them represent the parameters entering the FE-CT model Hamiltonian for excited states. Excitonic models, properly extended to include the coupling to quantum or classical vibrations find also application in the description of steady-state optical processes62 and real-time dynamics.64, 63, 65
@@
The quality of first-principle calculations can be preserved to a large extent by computing the model parameters ab initio in the spirit of a multiscale approach.62, 63, 65 For instance, XT couplings between singlet excitons are the only parameters in play in the standard FE model, and several methods have been proposed to its calculations at various levels of theory.66 The simplest approach is the so-called dimer splitting method that, however, strictly applies only to pairs of molecules equivalent by symmetry, each of which presents excitons well-separated in energy. A common approximation consists of neglecting the contributions due to the inter-molecular overlap (exchange terms, e.g. Dexter) and retaining the Coulomb, Förster-type, interaction between the transition densities of FEs, which is the leading term at a large inter-molecular distance. Point-dipole and distributed-monopole67 approximations of the transition density offer practical schemes for computing this interaction.
@@
Whenever inter-molecular overlap is of concern, or for more complex FE-CT exciton models, the derivation of parameters from first-principles usually follows so-called diabatization schemes,68, 69, 70, 71, 64, 72, 73, 74, 75, 76, 77 in which a unitary transformation is defined transforming the diagonal adiabatic representation of the Hamiltonian of the multichromophoric system into the diabatic one, featuring diabatic-states energies on the diagonal and the couplings among them elsewhere. The choice of the unitary transformation is somewhat arbitrary. Common choices employ the eigenstates of a physically-motivated observable (e.g. in the popular fragment excitation difference scheme, FED70), or projective schemes where suitable basis functions are introduced.72, 64, 77 Projective schemes based on BSE/ calculations, within the Tamm-Dancoff approximation (TDA),78, 79 have been proposed by Wehner and Baumeier80 and Leng et al..81 schemes in a very recent work.82 Full and TDA BSE/ calculations have been very recently combined with common diabatization methods (Edmiston-Ruedenberg,83 generalized Mulliken-Hush,68 fragment charge difference84) by Tirimbó and Baumeier.82
@@ 2 Theory
The present development aims to set up effective excitonic models within the framework of the many-body ab initio Green’s function techniques and BSE. This subject is briefly outlined here from a practical perspective, highlighting the features that will be mostly relevant to this end. For further detail, we refer the readers to review papers14, 31, 15, 16 and to the original publications of our implementation.23, 24, 44, 48, 49
Methods based on Green’s functions rely on many-body perturbation theory to compute electronic excitations upon introducing electronic correlation effects on top of mean-field electronic structure calculations, such as Kohn-Sham DFT. 85 The set of Kohn-Sham orbitals and one-particle energies obtained with a given functional approximation define the non-interacting system used as a starting point for the perturbation theory. Then, within this framework, charged excitations such as electron removal (e.g. ionization potential) and electron addition (e.g. electron affinity) can be computed as quasiparticle energies
where, in practice, the self energy replaces the Kohn-Sham DFT exchange-correlation potential . The self-energy is a non-local and energy-dependent operator, . This, within the formalism, is approximated as the product of the single-particle Green’s function () and the screened Coulomb interaction (), i.e. with i being the imaginary unit. In the so-called scheme, is built directly with the Kohn-Sham levels. Re-computing the self-energy with quasiparticle energies iteratively until convergence leads to the partially self-consistent ev scheme. This approach largely reduces the dependence of quasiparticle excitations on the starting DFT functional, leading to a good agreement with experimental values or high-level CCSD(T) calculations.32, 40, 86
Neutral excitations can be obtained via the BSE, which is a two-body correlated electron-hole (e-h) problem described by the following non-Hermitian eigenvalue equation
@@
where is the bare Coulomb interaction and is the screened one in the static limit (). Let us comment that the Kohn-Sham orbitals are real-valued in the present work. The anti-resonant blocks and are neglected in the Tamm-Dancoff approximation 78, 79, which is not adopted in this work. We emphasize that the screened-Coulomb potential matrix elements in the resonant block do not cancel when occupied states do not overlap with unoccupied states. This explains the success of BSE for CT excitations.
@@
We recall that, despite the similarity between equation 3 and the popular Casida’s linear-response equation in TD-DFT,5 the explicit account for e-h interaction permits a faithful description of excitations, regardless of their nature (FE or CT) or of the Kohn-Sham-DFT functional approximation employed to construct the single-particle basis. This determines an important qualitative and quantitative improvement over TD-DFT, especially for the stability of CT excitation energies, making BSE a robust and predictive tool for describing FE and CT excitons on an equal footing.87
@@ 2.2 Environmental embedding in /BSE
Many-body calculations of molecules and small molecular clusters in condensed phases have been performed by adopting hybrid quantum mechanics/classical mechanics (QM/MM) embedding techniques.44, 48, 49, 31 These approaches require accounting for the effect of the embedding environment at different steps of the computational workflow, first at the ground-state Kohn-Sham DFT level and then in the calculation of charged and neutral excitations.
@@
Environmental embedding in the ground state is essential for molecules dissolved in polar solvents, as well as in the solid state. Ground-state electrostatic embedding is indeed necessary to describe the strong crystal fields sourced by molecular multipole moments. A proper description of internal fields is essential for the quantitative description of the absolute binding energies in photoemission experiments on molecular solids,49 as well as to capture the dependence of energy levels on molecular orientations at film and crystal surfaces.56 In the following, ground-state embedding will be either based on PCM,44 to describe a generic non-polar medium, or on the charge response polarizable atomistic model,88, 89 whenever molecular crystals are the explicit target.
@@
The treatment of the screening of charged () and neutral (BSE) excitations in the presence of an environment has been documented in our original works.44, 48, 49 The key for this development consists in treating the dielectric susceptibility of the embedding medium as a frequency-independent quantity, i.e. the environment reacting instantaneously to the excitations within the QM region. The frequency-independent response of the environment is fully consistent with the BSE, which also relies on a static screening of the e-h interaction (see Eq. 4). The susceptibility is instead an explicitly dynamic quantity in the formalism, requiring an additional step for combining the dielectric response of QM and MM regions. The effect of the screening of charged excitations (polarization energy) can be computed at the static COHSEX level, i.e. the frequency-independent analog of .11
@@
In previous works, the state-specific polarization energy was computed as the difference between the quasiparticle energy obtained in two independent COHSEX calculations including or not the effect of the environment. Then, this difference was added as a perturbative correction to quasiparticle energies (COHSEX).48, 49 An alternative treatment of the environment emerged in the context of a QM/QM’ framework based on a fragment formalism.51 This approach allows an explicit account for the dynamics of the embedding medium (frequency ), but also to recover the instantaneous dielectric response of the environment in the limit. This permits the description of the dielectric embedding beyond the perturbative COHSEX approach. The novel self-consistent COHSEX treatment of environmental screening has been combined with PCM and adopted in the present work. The details of the methodology will be disclosed in a forthcoming publication.
@@ 2.3 Projective diabatization method
We consider a system composed of distinct molecular fragments, where we perform BSE calculations on the single fragments and for the whole system. For the sake of simplicity, a system composed of two fragments (dimer) is considered in the following. The generalization to ensembles of more fragments is straightforward. We define a set of e-h states that will be used as a diabatic basis to set up the exciton-model Hamiltonian, including FE excitons localized on one of the two fragments and inter-fragment CT states. Diabatic FE states are set to the BSE eigenstates of individual fragments, with corresponding to the th exciton of fragment . CT basis states are defined as with referring to an occupied Kohn-Sham orbital of fragment and to a virtual state of fragment . The number of selected FE and CT basis states (problem dimension , fixed by the range spanned by , and ) can be determined a posteriori, based on the energy range the model is expected to cover.
@@
This expression directly follows from the application to of the identity operator, in the form of a completeness relation for a non-orthogonal basis. The diagonal matrix , with elements , ensures the normalization of the target functions, but not their orthogonality. We insist on the fact that while the BSE wavefunctions of the dimer () form an orthonormal set, the target eigenstates for the model Hamiltonian () do not. Starting from such a common background, two methods can be proposed to obtain the Hamiltonian.
@@
where and Tr denotes the trace of the matrix . Then, expressing Eq. 14 in terms of traces of product of matrices and following the usual rules for the derivatives of matrices 90, we arrive at
that corresponds to a Sylvester’s equation 91, which allows us to find the Hamiltonian that is symmetric and best reproduces BSE eigenvalues and the target functions. Equation 15 can be solved with the Bartels–Stewart algorithm,92 see SI for details.
2.4 Approximate FE-FE and FE-CT exciton couplings
Besides the projective diabatization scheme described above, we also compute exciton couplings within common approximations. In the limit of zero intermolecular overlap, the coupling between singlet FE excitons reduces to the electrostatic interaction between transition densities.61 For a pair of molecules in the vacuum the FE-FE coupling is

### v2

@@ 1 Introduction
The accurate first-principles description of elementary excitations (excitons) in organic materials represents an invaluable tool for understanding their photophysical properties as well as phenomena underlying applications in optoelectronics, such as energy transfer, exciton splitting or charge separation in organic solar cells.1, 2, 3, 4
@@
Time-dependent density-functional theory (TD-DFT), in its linear-response formulation within the adiabatic approximation,5 has been the most popular method for the characterization of excitons in molecular systems. TD-DFT arguably represents one of the most reliable methods among those being computationally affordable for large and complex real-life materials. However, TD-DFT is known to present some limitations, such as the inaccurate and strongly functional-dependent description of CT excitations. This results in a possibly incorrect state ordering between CT and FE excitons,6, 7 or in missing the characteristic Coulomb-like dependence of CT states energy with inter-molecular distance when using local, semilocal, or global hybrid functionals.8 These shortcomings can be significantly attenuated with the use of range-separated functionals,9, 10 leading to results that however depend on the the specific parameterization for the range separation of the Coulomb repulsion term. 11, 12, 13
@@
Many-body perturbation methods based on Green’s function theory, such as the 14, 15, 16, 17, 18, 19 and the Bethe-Salpeter equation (BSE) formalisms,20, 21, 22, 23 stand as robust and affordable options for computing the excited states in molecular systems.24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34 Within the resolution-of-the-identity (RI-V) framework,35 and for converging iteratively a few low-lying BSE eigenstates, our /BSE implementation scales as . Besides, /BSE provides an accurate description of excitations of different natures thanks to a proper treatment of electronic correlations. Benchmark studies against reference methods and experimental data demonstrated a typical accuracy of 0.1-0.2 eV for charged excitations for valence states and optical transitions, when a self-consistent calculation of quasiparticle energies is performed.36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47 The explicit accounting for the screened non-local electron-hole (e-h) interaction permits a faithful description of excitations regardless of their nature (FE vs. CT) and of the type of system under scrutiny (e.g. organic homo- and hetero-molecular, inorganic, hybrid). Importantly, the use of a self-consistent scheme largely reduces the dependence of charged and neutral excitations on the functional used as starting point for the many-body perturbation theory.34, 38, 40, 45 We note that a residual starting-point dependence has been reported for neutral excitations, the magnitude of this fluctuation typically being comparable to the method’s accuracy.48, 49
@@
The possibility of describing molecular excitations in condensed phases with Green’s function many-body method has been enabled by the recent development of a multiscale embedding method of different levels of detail, namely polarizable continuum model (PCM),50, 51, 52, 53 classical atomistic (polarizable QM/MM)54, 55, 33 and full-quantum (QM/QM’)56, 57, 58. Embedded and BSE formalisms have been successfully applied in the contexts of molecular doping59, 60, 61, 62 and organic photovoltaics.63 Many-body perturbation theory methods have been historically developed for extended (i.e. periodic) systems17 and applications to molecular solids have been reported in the literature.64, 65, 66, 67 These calculations are extremely expensive and hardly affordable for realistic molecular solids encompassing many atoms in the unit cell or featuring structural disorder.
@@
Effective model Hamiltonians for excited states represent a cheap and insightful alternative to a full ab initio treatment for the description of the low-energy physics of multichromophoric systems.68, 3, 69, 70 These approaches rely on the definition of a reduced set of diabatic states characterized by wave functions localized on single molecular units, usually including pure intra-molecular FE and pure inter-molecular CT excitations with electron and hole localized on different fragments (see Figure 1). For this reason, these approaches are sometimes referred to as “low-energy” or “few-state” models. A typical model Hamiltonian for a molecular solid can be expressed as the sum of three terms,
@@
describing FEs (standard Frenkel exciton model), inter-molecular CT states, and the interaction between the two types of exciton (see SI for explicit expressions). Black arrows in Fig. 1 show the elementary processes corresponding to the quantum coupling between states, namely FE transfer(XT), excited-state hole/electron transfer (XHT/XET), and hole/electron transfer (HT/ET). The energy of the diabatic basis states and the coupling between them represent the parameters entering the FE-CT model Hamiltonian for excited states. Excitonic models, properly extended to include the coupling to quantum or classical vibrations find also application in the description of steady-state optical processes69 and real-time dynamics.71, 70, 72
@@
The quality of first-principle calculations can be preserved to a large extent by computing the model parameters ab initio in the spirit of a multiscale approach.69, 70, 72 For instance, XT couplings between singlet excitons are the only parameters in play in the standard FE model, and several methods have been proposed to its calculations at various levels of theory.73 The simplest approach is the so-called dimer splitting method that, however, strictly applies only to pairs of molecules equivalent by symmetry, each of which presents excitons well-separated in energy. A common approximation consists of neglecting the contributions due to the inter-molecular overlap (exchange terms, e.g. Dexter) and retaining the Coulomb, Förster-type, interaction between the transition densities of FEs, which is the leading term at a large inter-molecular distance. Point-dipole and distributed-monopole74 approximations of the transition density offer practical schemes for computing this interaction.
@@
Whenever inter-molecular overlap is of concern, or for more complex FE-CT exciton models, the derivation of parameters from first-principles usually follows so-called diabatization schemes,75, 76, 77, 78, 71, 79, 80, 81, 82, 83, 84 in which a unitary transformation is defined transforming the diagonal adiabatic representation of the Hamiltonian of the multichromophoric system into the diabatic one, featuring diabatic-states energies on the diagonal and the couplings among them elsewhere. The choice of the unitary transformation is somewhat arbitrary. Common choices employ the eigenstates of a physically-motivated observable (e.g. in the popular fragment excitation difference scheme, FED77), or projective schemes where suitable basis functions are introduced.79, 71, 84 Projective schemes based on BSE/ calculations, within the Tamm-Dancoff approximation (TDA),85, 86 have been proposed by Wehner and Baumeier87, 88 and Leng et al..89 in recent works. Also, full and TDA BSE/ calculations have been recently combined with common diabatization methods (Edmiston-Ruedenberg,90 generalized Mulliken-Hush,75 fragment charge difference91) by Baumeier and coworkers.88
@@ 2 Theory
The present development aims to set up effective excitonic models within the framework of the many-body ab initio Green’s function techniques and BSE. This subject is briefly outlined here from a practical perspective, highlighting the features that will be mostly relevant to this end. For further detail, we refer the readers to review papers17, 34, 18, 19 and to the original publications of our implementation.26, 27, 50, 54, 55
Methods based on Green’s functions rely on many-body perturbation theory to compute electronic excitations upon introducing electronic correlation effects on top of mean-field electronic structure calculations, such as Kohn-Sham DFT. 92 The set of Kohn-Sham orbitals and one-particle energies obtained with a given density functional approximation define the non-interacting system used as a starting point for the perturbation theory. Then, within this framework, charged excitations such as electron removal (e.g. ionization potential) and electron addition (e.g. electron affinity) can be computed as quasiparticle energies
where, in practice, the self energy replaces the Kohn-Sham DFT exchange-correlation potential . The self-energy is a non-local and energy-dependent operator, . This, within the formalism, is approximated as the product of the single-particle Green’s function () and the screened Coulomb interaction (), i.e. with i being the imaginary unit. In the so-called scheme, is built directly with the Kohn-Sham levels. Re-computing the self-energy with quasiparticle energies iteratively until convergence leads to the partially self-consistent ev scheme. This approach largely reduces the dependence of quasiparticle excitations on the starting DFT functional, leading to a good agreement with experimental values or high-level CCSD(T) calculations.36, 44, 93
Neutral excitations can be obtained via the BSE, which is a two-body correlated e-h problem described by the following non-Hermitian eigenvalue equation
@@
where is the bare Coulomb interaction and is the screened one in the static limit (). Let us comment that the Kohn-Sham orbitals are real-valued in the present work. The anti-resonant blocks and are neglected in the Tamm-Dancoff approximation 85, 86, which is not adopted in this work. We emphasize that the screened-Coulomb potential matrix elements in the resonant block do not cancel when occupied states do not overlap with unoccupied states. This explains the success of BSE for CT excitations.
@@
We recall that, despite the similarity between equation 3 and the popular Casida’s linear-response equation in TD-DFT,5 the explicit account for the screened e-h interaction permits a faithful description of excitations, regardless of their nature (FE or CT) or of the Kohn-Sham-DFT density functional approximation employed to construct the single-particle basis. This determines an important qualitative and quantitative improvement over TD-DFT, especially for the stability of CT excitation energies, making BSE a robust and predictive tool for describing FE and CT excitons on an equal footing.94
@@ 2.2 Environmental embedding in /BSE
Many-body calculations of molecules and small molecular clusters in condensed phases have been performed by adopting hybrid quantum mechanics/classical mechanics (QM/MM) embedding techniques.50, 54, 55, 34 These approaches require accounting for the effect of the embedding environment at different steps of the computational workflow, first at the ground-state Kohn-Sham DFT level and then in the calculation of charged and neutral excitations.
@@
Environmental embedding in the ground state is essential for molecules dissolved in polar solvents, as well as in the solid state. Ground-state electrostatic embedding is indeed necessary to describe the strong crystal fields sourced by molecular multipole moments. A proper description of internal fields is essential for the quantitative description of the absolute binding energies in photoemission experiments on molecular solids,55 as well as to capture the dependence of energy levels on molecular orientations at film and crystal surfaces.63 In the following, ground-state embedding will be either based on PCM,50 to describe a generic non-polar medium, or on the charge response polarizable atomistic model,95, 96 whenever molecular crystals are the explicit target.
@@
The treatment of the screening of charged () and neutral (BSE) excitations in the presence of an environment has been documented in our original works.50, 54, 55 The key for this development consists in treating the dielectric susceptibility of the embedding medium as a frequency-independent quantity, i.e. the environment reacting instantaneously to the excitations within the QM region. The frequency-independent response of the environment is fully consistent with the BSE, which also relies on a static screening of the e-h interaction (see Eq. 4). The susceptibility is instead an explicitly dynamic quantity in the formalism, requiring an additional step for combining the dielectric response of QM and MM regions. The effect of the screening of charged excitations (polarization energy) can be computed at the static COHSEX level, i.e. the frequency-independent analog of .14
@@
In previous works, the state-specific polarization energy was computed as the difference between the quasiparticle energy obtained in two independent COHSEX calculations including or not the effect of the environment. Then, this difference was added as a perturbative correction to quasiparticle energies (COHSEX).54, 55 An alternative treatment of the environment emerged in the context of a QM/QM’ framework based on a fragment formalism.57 This approach allows an explicit account for the dynamics of the embedding medium (frequency ), but also to recover the instantaneous dielectric response of the environment in the limit. This permits the description of the dielectric embedding beyond the perturbative COHSEX approach. The novel self-consistent COHSEX treatment of environmental screening has been combined with PCM and adopted in the present work. The details of the methodology will be disclosed in a forthcoming publication.
@@ 2.3 Projective diabatization method
We consider a system composed of distinct molecular fragments, where we perform BSE calculations on the single fragments and for the whole system. For the sake of simplicity, a system composed of two fragments (dimer) is considered in the following. The generalization to ensembles of more fragments is straightforward. We define a set of e-h states that will be used as a diabatic basis to set up the exciton-model Hamiltonian, including FE excitons localized on one of the two fragments and inter-fragment CT states. Diabatic FE states are set to the BSE eigenstates of individual fragments, with corresponding to the th exciton of fragment . CT basis states are defined as with referring to an occupied Kohn-Sham orbital of fragment and to a virtual state of fragment . The number of selected FE and CT basis states (problem dimension , fixed by the range spanned by , and ) can be determined a posteriori, based on the energy range the model is expected to cover (see below).
@@
This expression directly follows from the application to of the identity operator, in the form of a completeness relation for a non-orthogonal basis. The diagonal matrix , with elements , ensures the normalization of the target functions, but not their orthogonality. We note that while the BSE wavefunctions of the dimer () form an orthonormal set, the target eigenstates for the model Hamiltonian () do not. Starting from such a common background, two methods can be proposed to obtain the Hamiltonian.
@@
where and Tr denotes the trace of the matrix . Then, expressing Eq. 14 in terms of traces of product of matrices and following the usual rules for the derivatives of matrices 97, we arrive at
that corresponds to a Sylvester’s equation 98, which allows us to find the Hamiltonian that is symmetric and best reproduces BSE eigenvalues and the target functions. Equation 15 can be solved with the Bartels–Stewart algorithm,99 see SI for details.
Before closing this section, we point out some practical considerations regarding the selection of the number and nature of basis states.
The construction of a low-energy model relies on the somewhat arbitrary choice on the energy cutoff between the degrees of freedom to include in the model and those to discard.
This choice must be physically motivated, it should reduce as much as possible the dimension of the problem, but also
take into account common sense arguments.
For instance, the energy gap between the highest-energy target BSE exciton and the first-excluded one should be sizeable.
We advise against cutting through manifolds of degenerate or quasi-degenerate states.
The number of basis states must then equal the number of target states.
The matrix in Eq. 10 can be first evaluated for a larger set of candidate states, possibly including multiple FE and CT excitons per fragment.
The inspection of permits to identify the states that would finally constitute the basis, i.e. the state for which the projection of the dimer BSE excitons is the largest.
For example, the minimal basis for molecular homodimers consists in one FE per fragment, albeit in some cases the inclusion of low-lying CT states might be appropriate, as in the pentacene case study developed below.
In molecular heterodimers, such as donor-acceptor complexes, the inclusion in the basis of CT transitions from the occupied levels of the donor to the unoccupied subspace of the acceptor is crucial.
2.4 Approximate excitonic couplings
Besides the projective diabatization scheme described above, we also compute exciton couplings within common approximations. The FE-CT and some of the CT-CT couplings can be approximated as the matrix elements of the Kohn-Sham Hamiltonian between frontier molecular orbitals of the involved fragments. Referring to the example in Figure 1 for two-level molecules (HOMO, LUMO), the one-electron approximation for the coupling associated with XHT and HT process between fragments and is , where is the Kohn-Sham Hamiltonian for the dimer - and is HOMO of fragment . Similarly, the matrix element for XET and ET processes reads .
In the limit of zero intermolecular overlap, the coupling between singlet FE excitons reduces to the electrostatic interaction between transition densities.68 For a pair of molecules in the vacuum the FE-FE coupling is